OpenStudy (anonymous):

to estimate the height of building, a stone is dropped from the top of the building into pool of water at ground level. how high is the building if the splash is seen 5.6 seconds after the stone is dropped

OpenStudy (akashdeepdeb):

Hello @McCoolmisfit . :) Is this a physics question or a math question? Time to reach the ground: 5.6 s Height of building: x meters. Acceleration due to gravity: 9.8 m/s\(^2\) Initial velocity: 0m/s [Initially the stone must have been dropped with no velocity] Final Velocity: ? Now you can use some laws of Physics. h = ut + 1/2 gt^2 = 0 + 1/2 * 9.8 * (5.6)^2 = 4.9 * 5.6 * 5.6 m This must be the height of the building. :) Getting this?